Problem: x^2 +8x +52 =0
Step-by-step, where do I go from here? I really need this spelled out for me so that I can actually understand it. I don't just want the answer, I want to understand how it is worked. :-)
Thanks!
Angela

Answer by stanbon(75887)

x^2 +8x +52 =0
---
Quadratic Formula:
x = [-b +- sqrt(b^1-4ac)]/(2a)
-----
x = [-8 +- sqrt(64-4*1*52)]/2
-----
x = [-8 +- sqrt(-144)]/2
-----
x = [-8 +- 12i]/2
---
x = (-8/2) +- 12i/2
---------
x = -4 +- 6i
-------
x = -4+6i or x = -4-6i
===================
Cheers,
Stan H.
===================